OpenStudy (kendricklamar2014):

Determine if the number is written in scientific notation. if not, explain 32 * 10^4. A. No; it is not written as a power of 10 B. No; the first factor is not a number between 1 and 10 C. Yes; the number is written in scientific notation

OpenStudy (anonymous):

B

OpenStudy (solomonzelman):

a number written in scientific notation has to be: \(\large\color{black}{ \displaystyle {\rm number}~\times ~10^{{\rm n}} }\) where, 1. \(\large\color{black}{ \displaystyle 0> {\rm number}>10}\) 2. \(\large\color{black}{ {\rm n}~~~~~is~~an~~integer. }\)

OpenStudy (solomonzelman):

so this "number" (I will refer to it, as the "base") has to be between 0 and 10 as indicated, for this to be a properly written scientific notation. is 32 between 0 and 10?

OpenStudy (solomonzelman):

I didn't say that

OpenStudy (solomonzelman):

n (in the exponent) has to be an integer. "number" can be a non integer.

OpenStudy (kendricklamar2014):

No, 32 is not between 0 and 10

OpenStudy (anonymous):

Yes, I know. My bad :)

OpenStudy (solomonzelman):

Yes, exactly, so what is YOUR answer?

OpenStudy (kendricklamar2014):

B

OpenStudy (solomonzelman):

yes, it is correct.

OpenStudy (solomonzelman):

it is multiplied times 10^(some power) so that requirement it does pass...

OpenStudy (solomonzelman):

For optional... can you tell me how you would write it properly in order for it to be a scientific notation?

OpenStudy (solomonzelman):

if you don't know, you can just go ahead and say so...

OpenStudy (kendricklamar2014):

For it to be in Scientific Notation, the 32 would have to be a number between 0-10

OpenStudy (solomonzelman):

yes, and you know that: \(\large\color{black}{ \displaystyle 32\times 10^4 }\) is \(\large\color{black}{ \displaystyle 320,000}\) So, \(\large\color{black}{ \displaystyle 32\times 10^4 }\) would be same as \(\large\color{black}{ \displaystyle 3.2\times 10^5 }\) you are multiplying times 10 by the \(\large\color{black}{ \displaystyle 10^4 }\) and dividing by 10 by \(\large\color{black}{ \displaystyle 3.2 }\).

OpenStudy (solomonzelman):

So you aren't changing the value by dividing by 10 and multiplying by 10.

OpenStudy (solomonzelman):

and 3.2 satisfies the statement 0>3.2>10

OpenStudy (kendricklamar2014):

Oh, now I get it, thanks

OpenStudy (solomonzelman):

good. You wlecome!